OpenStudy (anonymous):

the problem i am having trouble with is... 9 x 3 + 2 x -2 (3--2)

OpenStudy (shadowfiend):

Ok, so first off let's get rid of that nasty double-negative. Subtracting a negative is the same as adding. So 3 - -2 is the same as 3 + 2. Given that, can you figure out the rest?

OpenStudy (anonymous):

9*3+2*-2(3--2) = (9*3)+2{-2(3--2)} first, 9*3=27 , and 3--2=3-(-2)=3+2=5 so, 9*3+2*-2(3--2) =27+2*(-2)*5=27+(-20)=7